If you have a spinning wheel with 3 green sectors, 1 blue sector and 1 red sector, what is the probability of getting a green sector? What is the probability of getting a non-blue sector?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

A total of five sectors are present.
Out of the five sectors, three sectors are green.
∴, probability of getting agreen sector = 3/5

Out of the five sectors, one sector is blue.
Hence, Non-blue sectors = 5–1 = 4 sectors
∴, probability of getting anon-blue sector= 4/5

Thanks